Calculate the number of moles in a sulphur atom..................
 
 

Grade:11

1 Answers

Arun
5 years ago

As we know that atomic weight of sulphur is 32i.e. according to the mole concept 32gram of sulphur contains 6.022×10^23 sulphur atoms

Hence using unitary method 1 gram of sulphur contain “6.022×10^23÷32″ atoms

Again then for no of sulphur atoms in 8 gram of sulphur equals to (6.022×10^23÷32 )×8 = 6.022×10^23÷4 = 1.5055×1023

This works on concept that “atomic weight of atom = 1 mole of that atom = 6.022×10^23 no of atoms”

Here we took “sulphur” so using this formule

32gram sulphur = 1 mole of sulphur=6.022×10^23sulphur atom